APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Friday January 1, 2010 (MJD 55198) ---=== OBSERVING TEAM ===--- Audrey Simmons Dan Oravetz Steph Snedden (afternoon) Dan Long (support) French Leger (support) ---=== OBSERVING PLAN ===--- MARVELS Science ---=== OBSERVING SUMMARY ===--- It looks like the DIMM needs to be focused. We noticed that the 2.5m wind speed was reading -35.9 mph during the night on the APO Weather Status page. Also, the dust counter was reading 0 a lot. We discovered that cart. 8 has the front black panel corresponding to cart. 6 (labeled as "ET cart. 6"). We noticed this after we mounted 6 on the telescope to end the night. All of the B-pointings had the wrong set of guide fibers X-ed out in the guider display. We had a problem opening the enclosure at sunset with a tripped breaker, but resetting it brought things back to normal. We're still unsure as to the root cause, however. There were about half a dozen small telescope jumps during exposures. The pointing wasn't very good tonight; most of the time that we weren't exposing was spent trying to manually find stars. The fiducials for the current pointing model were taken about 7C colder than conditions we were observing in. We were unable to observer 3704A (cart. 7) because we found guide stars here and there, but when we finally had three stars, the guider was showing error vectors in opposite directions and naturally was in a stalemate. Observed the following: cart. 6, 3513B cart. 9, 3706B cart. 8, 3713A & 3713B cart. 4, 3219A & 3219B cart. 1, 3711A cart. 5, 3714A & B cart. 7, 3704B We found some fibers that need to be checked, they're described in the Log. ---=== OBSERVING LOG ===--- Afternoon: ---------- French, Dan, Wendell, Norm and Bob were here until mid-afternoon repairing the torn cable from last night. An excerpt from French's apo-daylog: "The power cable for the alt Wind Baffle drive motor broke away from its hanger - we suspect the cold got to the wire ties - and it fall far enough to be caught by az support ring and rip itself out of it junction box and short the cable to ground. Luckily the WB amp has a crowbar feature that effectively acted as a fuse and no other damage then a severed cable was sustained. We repaired the cable and rehung it from its hanger using SS clamps instead of wire ties. We ran the telescope through it paces and all clearances are good and the telescope is happy once more." They also changed the burned out flat-field lamp. Thanks guys! Saved me from dragging a ladder into a tight space. The rest of afternoon checkout was problem-free. 22:20Z - ThAr lamp turned on. Guider dark and flat taken. 20-second Spectro dark taken 23:32Z - Enclosure ventilation started, doors opened. Cartridge 6 mounted. Night: ------ 00:01Z - We had a problem opening the enclosure. After the enlcosure had run about a foot away from the stop, the power cut out and the high-voltage box on the S side of the upper enclosure was reading "Fault." We had checked for icicles on the power rails and didn't see anything prior to opening, but we checked again and still didn't see any sign of ice. We contacted Dan Long and French and both suggested checking the breaker in the lower enclosure on the NE side. We saw that the upper enclosure roll-switch was tripped, so we reset it and then the enclosure rolled off without a hiccup. 00:48Z - Calibrations. Swapped the spectro dewar after the fill completed. We wanted to be sure that there was enough LN2 for the morning since there was only one fill left in the dewar and we were anticipating a lot of plugging, so we thought we'd just get it done with. Cart. 6, 3513B, FIELD1572 ************************* 01:09Z - gotoField. Applied an offset manually (+0.002 in dec.) in order to acquire the field. 01:11Z - When the guider was turned on, the wrong set of guide fibers were X-ed out. We manually enabled and disabled fibers until the correct set of fibers were active. We later noticed this behavior for all B-pointings. 01:17Z - Science. 01:24Z - We saw a drop in flux on the PMT plot and upon subsequently checking the guider we saw a small telescope jump. See guider frames 0052-0053. 02:09Z - Calibrations. Cart. 9, 3706B, HD219828 ************************ 02:17Z - gotoField. Applied an offset manually (+0.004 in dec.) in order to acquire the field. 02:20Z - Science. 02:40Z - We noticed another small telescope jump, see guider frames 342-343. 03:11Z - Calibrations. The fiber corresponding to spectra (057,058) - blue green cyan, needs to be checked. Cart. 8, 3713A, HD4203 ********************** 03:17Z gotoField. Applied an offset manually (+0.004 in dec. and then +0.002 in RA) in order to acuire the field. 03:22Z - Science. 04:13Z - Calibrations. The fibers corresponding to spectra (047,048) - green black orange, (019,020) - green green cyan, and (007,008) - green black yellow, seem to be dropped or broken. Cart. 8, 3713B, HIP14810 ************************ 04:14Z - gotoField. Applied an offset manually (+0.002 in dec.) in order to acquire the field. 04:22Z - Science. 05:14Z - Calibrations. The fiber for spectra (003,004) - green black cyan, needs to be checked. Cart. 4, 3219A, HD17156 *********************** 05:19Z - gotoField. GSOGTF. 05:23Z - Science. 05:38Z - We noticed a small telescope jump (guider images 1156-58). 06:17Z - Calibrations. Cart. 4, 3219B, HD68988 *********************** 06:17Z - gotoField. It took us a long time to find guide stars. We just poked around with guesses on offsets since the offsets from last night's log didn't get us there. The magic combo was finally -0.003 in RA and -0.006 in dec. 06:54Z - Science. 07:45Z - Calibrations. Cart. 1, 3711A, HD80606 *********************** 07:50Z - gotoField. Used -0.004 offset in RA to acquire the field. 08:00Z - Science. 08:51Z - Calibrations. Cart. 7, 3704A, HD88133 *********************** 08:55Z - gotoField. Audrey manually did a spiral pattern in order to find stars. We eventually came across a star in fiber 5 that was extremely bright - plenty saturated with even 2 sec. No other stars were seen and the guider couldn't centroid well on the saturated star, so we tried other offsets. We found a star in fiber 11, but no other stars came into view while 11 was trying to center up. We went back to the bright star in fiber 5 and then played with rotator offsets and found a star in fiber 7. We finally ended up with stars in 1, 9, and 11, but the guider would not bring them in; in order to bring the stars in required telescope motions in different directions that would bring one in and push one out (see, for example, the error vectors in guider image 1984). 09:51Z - Calibrations. Cart. 5, 3714A, 47UMA ********************* 09:55Z - gotoField. Manually found the field by playing with offsets. We noticed another telescope jump, see guider image 2080. 10:15Z - Science. 11:06Z - Calibrations. Cart. 5, 3714B, HD118203 ************************ 11:06Z - gotoField. Applied same offset as field A in order to find the field. 11:15Z - Science. We saw another telescope jump, see guider frame 2431. 12:09Z - Calibrations. Cart. 7, 3704B, XO-1 ******************** 12:13Z - gotoField. Audrey used the same offsets as the previous plate since the pointing was similar and we managed to find stars. 12:18Z - Science. We saw another telescope jump at guider frame 2629. 13:11Z - Calibrations. Please check (115,116) - yellow clear cyan. 13:36Z - endNight rejoiced. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): y QA Procedures Done (y/n): y Name Exp. Time Flavor ============================================================================= MARVELS1.2010-01-02T00:48:17.340.fits 11.000000 TIO 41717.816406 MARVELS1.2010-01-02T00:50:24.010.fits 160.000000 THAR 252.449997 MARVELS1.2010-01-02T00:54:59.150.fits 11.000000 TIO 42625.183594 MARVELS1.2010-01-02T01:17:16.730.fits 3000.000000 STAR 31.245001 MARVELS1.2010-01-02T02:09:01.240.fits 11.000000 TIO 44471.089844 MARVELS1.2010-01-02T02:11:06.920.fits 160.000000 THAR 284.356262 MARVELS1.2010-01-02T02:15:43.240.fits 11.000000 TIO 44440.363281 MARVELS1.2010-01-02T02:20:01.110.fits 3000.000000 STAR 41.569668 MARVELS1.2010-01-02T03:11:45.110.fits 11.000000 TIO 45497.636719 MARVELS1.2010-01-02T03:13:51.260.fits 160.000000 THAR 279.393738 MARVELS1.2010-01-02T03:18:28.390.fits 11.000000 TIO 44562.273438 MARVELS1.2010-01-02T03:21:57.630.fits 3000.000000 STAR 31.406000 MARVELS1.2010-01-02T04:13:42.350.fits 11.000000 TIO 46608.453125 MARVELS1.2010-01-02T04:15:49.410.fits 160.000000 THAR 284.556244 MARVELS1.2010-01-02T04:20:25.670.fits 11.000000 TIO 45405.453125 MARVELS1.2010-01-02T04:22:23.920.fits 3000.000000 STAR 48.201668 MARVELS1.2010-01-02T05:14:09.760.fits 11.000000 TIO 45543.816406 MARVELS1.2010-01-02T05:16:16.350.fits 160.000000 THAR 280.899994 MARVELS1.2010-01-02T05:20:51.310.fits 11.000000 TIO 43410.546875 MARVELS1.2010-01-02T05:23:12.090.fits 3000.000000 STAR 48.246666 MARVELS1.2010-01-02T06:14:56.110.fits 11.000000 TIO 81.454544 MARVELS1.2010-01-02T06:17:02.020.fits 11.000000 TIO 42004.816406 MARVELS1.2010-01-02T06:19:08.090.fits 160.000000 THAR 258.718750 MARVELS1.2010-01-02T06:23:46.500.fits 11.000000 TIO 40341.273438 MARVELS1.2010-01-02T06:53:32.370.fits 3000.000000 STAR 34.757332 MARVELS1.2010-01-02T07:45:18.350.fits 11.000000 TIO 42849.183594 MARVELS1.2010-01-02T07:47:25.330.fits 160.000000 THAR 256.70623 MARVELS1.2010-01-02T07:52:00.490.fits 11.000000 TIO 42433.000000 MARVELS1.2010-01-02T07:59:57.410.fits 3000.000000 STAR 34.145668 MARVELS1.2010-01-02T08:51:56.980.fits 11.000000 TIO 42852.910156 MARVELS1.2010-01-02T08:54:05.360.fits 160.000000 THAR 267.825012 MARVELS1.2010-01-02T08:58:41.220.fits 11.000000 TIO 42185.363281 MARVELS1.2010-01-02T09:51:35.600.fits 11.000000 TIO 42394.183594 MARVELS1.2010-01-02T09:53:43.520.fits 160.000000 THAR 271.725006 MARVELS1.2010-01-02T09:58:20.620.fits 11.000000 TIO 42547.363281 MARVELS1.2010-01-02T10:15:04.160.fits 3000.000000 STAR 40.285000 MARVELS1.2010-01-02T11:06:50.090.fits 11.000000 TIO 40896.089844 MARVELS1.2010-01-02T11:08:56.560.fits 160.000000 THAR 263.899994 MARVELS1.2010-01-02T11:13:32.440.fits 11.000000 TIO 43742.726562 MARVELS1.2010-01-02T11:15:29.590.fits 3000.000000 STAR 42.557667 MARVELS1.2010-01-02T12:09:13.350.fits 11.000000 TIO 40977.363281 MARVELS1.2010-01-02T12:11:21.610.fits 160.000000 THAR 264.012512 MARVELS1.2010-01-02T12:15:57.770.fits 11.000000 TIO 43461.636719 MARVELS1.2010-01-02T12:18:50.270.fits 3000.000000 STAR 24.636667 MARVELS1.2010-01-02T13:11:27.800.fits 11.000000 TIO 38736.816406 MARVELS1.2010-01-02T13:13:35.020.fits 160.000000 THAR 250.606247 MARVELS1.2010-01-02T13:18:12.020.fits 11.000000 TIO 41978.453125 ---=== TELESCOPE OFFSETS AND SCALE I ===--- Cart. 6, 3513B -------------- ObjArcOff = 0.002093, 0.000000, 4769113855.13000, 0.002160, 0.000000, 4769113855.13000 GuideOff = 0.000000, 0.000000, 4769113855.13000, 0.000000, 0.000000, 4769113855.13000, 0.024494, 0.000000, 4769113855.13000 CalibOff = 0.000271, 0.000000, 4769113855.13000, 0.000248, 0.000000, 4769113855.13000, 0.000000, 0.000000, 4769113855.13000 Cart. 9, 3706B -------------- ObjArcOff = 0.001900, 0.000000, 4769115847.62000, 0.003910, 0.000000, 4769115847.62000 GuideOff = 0.000000, 0.000000, 4769115847.62000, 0.000000, 0.000000, 4769115847.62000, 0.014800, 0.000000, 4769115847.62000 CalibOff = 0.000271, 0.000000, 4769115847.62000, 0.000248, 0.000000, 4769115847.62000, 0.000000, 0.000000, 4769115847.62000 Cart. 8, 3713A -------------- ObjArcOff = 0.002438, 0.000000, 4769119495.79000, 0.003905, 0.000000, 4769119495.79000 GuideOff = 0.000000, 0.000000, 4769119495.79000, 0.000000, 0.000000, 4769119495.79000, 0.020996, 0.000000, 4769119495.79000 CalibOff = 0.000271, 0.000000, 4769119495.79000, 0.000248, 0.000000, 4769119495.79000, 0.000000, 0.000000, 4769119495.79000 Cart. 8, 3713B -------------- ObjArcOff = 0.001066, 0.000000, 4769123386.37000, 0.004058, 0.000000, 4769123386.37000 GuideOff = 0.000000, 0.000000, 4769123386.37000, 0.000000, 0.000000, 4769123386.37000, 0.037115, 0.000000, 4769123386.37000 CalibOff = 0.000271, 0.000000, 4769123386.37000, 0.000248, 0.000000, 4769123386.37000, 0.000000, 0.000000, 4769123386.37000 Cart. 4, 3219A -------------- ObjArcOff = -0.001510, 0.000000, 4769127174.95000, 0.000081, 0.000000, 4769127174.95000 GuideOff = 0.000000, 0.000000, 4769127174.95000, 0.000000, 0.000000, 4769127174.95000, 0.027648, 0.000000, 4769127174.95000 CalibOff = 0.000271, 0.000000, 4769127174.95000, 0.000248, 0.000000, 4769127174.95000, 0.000000, 0.000000, 4769127174.95000 Cart. 4, 3219B -------------- ObjArcOff = -0.003345, 0.000000, 4769134414.59000, 0.001123, 0.000000, 4769134414.59000 GuideOff = 0.000000, 0.000000, 4769134414.59000, 0.000000, 0.000000, 4769134414.59000, 0.006168, 0.000000, 4769134414.59000 CalibOff = 0.000000, 0.000000, 4769134414.59000, 0.000000, 0.000000, 4769134414.59000, 0.000000, 0.000000, 4769134414.59000 Cart. 1, 3711A -------------- 0ObjArcOff = -0.004625, 0.000000, 4769136347.29000, 0.000935, 0.000000, 4769136347.29000 GuideOff = 0.000000, 0.000000, 4769136347.29000, 0.000000, 0.000000, 4769136347.29000, 0.023348, 0.000000, 4769136347.29000 CalibOff = 0.000000, 0.000000, 4769136347.29000, 0.000000, 0.000000, 4769136347.29000, 0.000000, 0.000000, 4769136347.29000 Cart. 5, 3714A -------------- ObjArcOff = -0.003278, 0.000000, 4769146907.36000, 0.000620, 0.000000, 4769146907.36000 GuideOff = 0.000000, 0.000000, 4769146907.36000, 0.000000, 0.000000, 4769146907.36000, -0.000411, 0.000000, 4769146907.36000 CalibOff = 0.000000, 0.000000, 4769146907.36000, 0.000000, 0.000000, 4769146907.36000, 0.000000, 0.000000, 4769146907.36000 Cart. 5, 3714B -------------- ObjArcOff = -0.003286, 0.000000, 4769150435.13000, 0.002297, 0.000000, 4769150435.13000 GuideOff = 0.000000, 0.000000, 4769150435.13000, 0.000000, 0.000000, 4769150435.13000, 0.002983, 0.000000, 4769150435.13000 CalibOff = 0.000000, 0.000000, 4769150435.13000, 0.000000, 0.000000, 4769150435.13000, 0.000000, 0.000000, 4769150435.13000 Cart. 7, 3704B -------------- ObjArcOff = -0.003669, 0.000000, 4769152385.06000, 0.005255, 0.000000, 4769152385.06000 GuideOff = 0.000000, 0.000000, 4769152385.06000, 0.000000, 0.000000, 4769152385.06000, 0.025443, 0.000000, 4769152385.06000 CalibOff = 0.000000, 0.000000, 4769152385.06000, 0.000000, 0.000000, 4769152385.06000, 0.000000, 0.000000, 4769152385.06000 ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ 01:48Z 6 3513B 1.00003 -296 1386 425 224 42.6 3.5 7 84 gdr. 1.7" 02:22Z 9 3706B 1.00009 -889 869 425 280 50.4 5.2 1 222 gdr. 1.7" 03:23Z 8 3713A 1.00017 -1678 172 400 282 56.2 4.6 1 258 gdr. 1.7" 04:28Z 8 3713B 1.00008 -790 824 350 308 72.3 4.4 1 219 gdr. 1.7" 05:31Z 4 3219A 0.99993 691 2089 325 196 46.7 3.8 1 294 gdr. 1.7" 07:34Z 4 3219B 0.99997 296 1698 300 165 59.8 1.7 2 234 gdr. 1.7" 08:06Z 1 3711A 1.00021 -2073 -251 325 143 65.4 1.0 3 304 gdr. 1.7" 11:02Z 5 3714A 1.00013 -1284 454 400 162 82.0 2.2 6 314 gdr. 1.7" 12:00Z 5 3714B 1.00015 -1480 273 350 145 61.6 2.4 8 347 gdr. 1.7" 12:33Z 7 3704B 1.00021 -2073 -157 350 100 42.0 1.7 10 1 gdr. 1.7" ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 22:12Z 43 3 0 264 (W) 0 - clear 22:45Z 42 4 0 334 (NNW) 0 - clear 23:19Z 43 0 3 352 (NNW) 0 - clear 23:52Z 42 -3 4 359 (NNW) 0 - clear 01:07Z 39 0 10 78 (ENE) 0 - clear 01:40Z 39 -2 6 86 (E) 0 - clear 02:13Z 40 -3 0 186 (S) 0 - clear 02:46Z 41 -2 1 222 (SW) 0 - clear 03:19Z 40 -1 1 244 (WSW) 0 - clear 03:52Z 39 -2 2 257 (WSW) 0 - clear 04:25Z 40 -1 1 216 (SW) 0 - clear 04:58Z 39 -3 1 248 (WSW) 0 - clear 05:31Z 38 0 1 293 (WNW) 0 - clear 06:04Z 37 5 1 286 (WNW) 0 - clear 06:38Z 36 5 1 208 (SSW) 0 - clear 07:11Z 36 4 1 241 (WSW) 0 - clear 07:44Z 35 11 2 282 (WNW) 0 - clear 08:17Z 36 6 6 275 (W) 0 - clear 08:50Z 36 1 9 314 (NW) 0 - clear 09:23Z 35 0 10 325 (NW) 0 - clear 09:56Z 34 1 2 342 (NNW) 0 - clear 10:29Z 34 4 9 327 (NNW) 0 - clear 11:02Z 35 1 6 321 (NW) 0 - clear 11:35Z 35 14 7 327 (NNW) 0 - clear 12:08Z 36 14 9 0 (N) 0 - clear ---=== TELESCOPE STATUS ===--- 23:32Z - Began ventilation. Doors and louvers opened, fans on. 00:20Z - Enclosure off. 13:28Z - Enclosure on, doors and louvers closed, fans off. No interlocks bypassed. Status at: 13:31Z Telescope stowed at: 121,30,00 mount Instrument mounted: Cartridge 6 Counterweights at: 243 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 275.70 lbs, 17.48 psi Imager N/A lbs, N/A psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_40_0 Watcher: v2_36_0 MCP: v6_6_1 TPM: tpm_v3_5_0 AstroDa: v15_13_1 TCC: TCC 2.11.0 2009-06-24 sdssProcedures: v3_02 SoS: v5_1_8 plate-mapper: v4_3_1 boss : version="v2_0_17" guider : guiderVersion="v1_0_16" hub : version="Tag: sdss3_v1_1_3" platedb : version="trunk+svn105808" sop : version="v1_0_4" ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A -0.7020 -0.6940 Axial B -0.0260 -0.0410 Axial C -0.3770 -0.3710 Trans D 11.6150 11.6170 Lateral E 10.3759 10.3505 Lateral F 11.9507 11.9253 GALILS Commanded: 9900. 400. 5850. 1200. 2950. 2750. Actual: 9917. 401. 5832. 1205. 2954. 2757. SETMIR VALUES PriDesOrient: 0.00 -3.30 5.10 278.08 178.06 PriOrient: 0.00 -3.04 5.26 277.70 178.07 SECONDARY: ---------- Focus: 0.00 Air Temp.: 4.9 Alt.: 29.998363 MIGS TONIGHT NOMINAL Axial A 18.6180 18.6150 Axial B 15.6220 15.6120 Axial C 13.8200 13.8170 Trans D -11.7720 -11.7780 GALILS Commanded: 1596713. 1567435. 1589696. -7600. -12400. Actual: 1597721. 1566188. 1588095. -7597. -12380. SETMIR VALUES SecDesOrient: 1257.00 -4.00 -9.00 10.00 234.42 SecOrient: 1257.49 -4.38 -8.22 10.70 234.94 ---=== PROBLEMS IN DETAIL ===---